OpenStudy (anonymous):

Solve by completing the square: x^2-5x+5=0

OpenStudy (anonymous):

Do you know how to complete the square?

OpenStudy (anonymous):

Yes. 1/2 the middle number then squared

OpenStudy (anonymous):

Ok, so complete this square.

OpenStudy (anonymous):

After getting the 5/2^2 which is 25/4, my question is what to do next.

OpenStudy (anonymous):

x^2 + bx + c = 0 -> [x^2 + bx + (b^2)/4] = (b^2)/4 - c (x + b/2)^2 = (b^2)/4 - c So, you will add that 25/4 to both sides and subtract 5 from both sides.

OpenStudy (anonymous):

So, at this point, you will have: (x - 5/2)^2 = 25/4 - 5

OpenStudy (anonymous):

Question #1: Do you understand the process up to this point? Question #2: can you take it from here? 2 questions. Looking for 2 answers.

OpenStudy (anonymous):

Is it (x-5/2)^2=5/4? or did I mess that up?

OpenStudy (anonymous):

You're doing fine. Now you have to get the square root. Don't forget the +- sign since you have a square root.

OpenStudy (anonymous):

x - 5/2 = +- sqrt(5/4)

OpenStudy (anonymous):

The answer in the book says 5+,-sqrt5/2 but what I ended up getting was 5/2+,-5/4

OpenStudy (anonymous):

x - 5/2 = +- sqrt(5/4) -> x - 5/2 = +- sqrt(5)/2 x = 5/2 +- sqrt(5)/2 -> x = [5 +- sqrt(5)] / 2

OpenStudy (anonymous):

mhmm.

OpenStudy (anonymous):

And that's it!

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

uw!